A=$875,P=$500,
r=?,t=30 years

Answer by stanbon(75887)

A=$875,P=$500,
r=?,t=30 years
-----------------------
Formula for simple interest:
A = Prt
---
875 = 500*r*30
875 = 15000r
r = 875/15000
r = 7/12 = 0.058333.. = 5.83%
=================================
Cheers,
Stan H.
====================